OpenStudy (anonymous):

Pure water has a pH level of 7.0. The pH scale is a measurement scale that indicates how acidic or basic (that is, alkaline) a liquid is. Each one-unit decrease in pH indicates a ten-fold increase in acidity. Similarly, for pH values greater than 7.0, each one-unit increase indicates an acidity level one-tenth of the next lower integer value. a) Black coffee has a pH level of 5. How much more acidic is black coffee than pure water? b) Ammonia has a pH of 11. What is its acidity compared to pure water? c) Lemon juice has a pH of 2. How does its acidity compare with milf of meagnesia

OpenStudy (anonymous):

a) black coffee = 7-5=2, therefore 10^2 times (100x) more acidic than pure water b) ammonia = 11-7= 4, therefore 10^4 times more basic (10000x) more basic